I am having a very hard time understanding the addition property of equality.
for axample 3x - 5 + 2x - 7 + x = 5x+2
The book says that the answer is 14, but how do you reach that point?

Answer by arunpaul(104)

3x -5 + 2x -7 + x = 5x +2
or (3x +2x +x -5x )= 2 +7 +5
or x = 14 (Ans) so simple